The two ends of a train moving with constant acceleration pass a certain point with velocities ‘u’ and ‘3u’.The velocity with which the middle point of the train passes the same point ??

Grade:

1 Answers

Arun
5 years ago
S = (v^2 - u^2) / (2a)
If Length of train is L then
L = ((3u)^2 - u^2) / (2a)
L = 4u^2 / a —-(1)
Midpoint is at Distance L/2
S = (v^2 - u^2) / (2a)
L/2 = (v^2 - u^2) / (2a)
From equation (1)
2u^2 / a = (v^2 - u^2) / (2a)
4u^2 = v^2 - u^2
v^2 = 5u^2
v = u√5
Velocity with which middle point passes the same point is u√5
 
 
Regards
Arun